# Optimal Investment Portfolio of Thyristor- and IGBT-based Electrolysis Rectifiers in Utility-scale Renewable P2H Systems

### v1

@@ IV-B Investment Portfolio Results of the Small-scale System for
For the small-scale system, with the intraday renewable power uncertainty set to zero, the rectifier configurations of four ELZs are evaluated. Fig. 10 presents the total revenues of different portfolios of TR and IGBT-R, while Table V summarizes the planning and operational results. Among the results, cases where IGBT-Rs are allowed to adjust reactive power are denoted ②-⑤, and cases where the PF of the IGBT-R is fixed at 1 are denoted ⑥-⑨ (see the discussion about reactive power characteristics in Section II-C-2).
As shown in Fig. 10, when the IGBT-Rs are allowed to provide reactive power, the optimal configuration is 3 TRs and 1 IGBT-R. In contrast, if the IGBT-R is fixed at , a uniform configuration of 4 TRs is preferable. The reasons are as follows.
@@
In terms of hydrogen production in configurations ① to ⑨, Table V shows that each additional IGBT-R reduces the hydrogen yield by approximately 2,500 kg/year, resulting in an estimated decrease in revenue of 4.75 CNY. On the other hand, the annual cost of an IGBT-R is 0.104(220–100) = 12.48 CNY greater than that of a TR considering the discount. Therefore, an IGBT-R investment is justified only if it reduces SVG capacity by at least (12.48 + 4.75)/(0.10435) = 4.73 MVar. In particular, Fig. 10 shows the SVG capacity with varying numbers of IGBT-Rs, along with the critical point at which configuring more IGBT-Rs is beneficial. It is clear that the only case ①② meets the conditions.
| ① | 4 + 0 | 2.0730 | 6.72 | 1362.64 | 1233.1 | ||||||||||
| ② | 3 + 1 | 2.0785 | 0.55 | 1360.27 | 1229.3 | ||||||||||
| ③ | 2 + 2 | 2.0641 | 0 | 1358.21 | 1218.0 | ||||||||||
| ④ | 1 + 3 | 2.0465 | 0 | 1355.46 | 1217.8 | ||||||||||
| ⑤ | 0 + 4 | 2.0289 | 0 | 1352.77 | 1211.8 | ||||||||||
| ⑥ | 3 + 1 | 2.0679 | 3.72 | 1360.77 | 1238.3 | ||||||||||
| ⑦ | 2 + 2 | 2.0618 | 0.56 | 1358.09 | 1235.5 | ||||||||||
| ⑧ | 1 + 3 | 2.0462 | 0 | 1355.32 | 1228.0 | ||||||||||
| ⑨ | 0 + 4 | 2.0288 | 0 | 1352.68 | 1217.9 |
@@ IV-C Operational Analysis of Different Rectifier Configurations
To analyze the operational differences of the ReP2H system under various rectifier configurations, we focus on configurations ①-③. Configurations ④ and ⑤ are similar to ③, whereas ⑥-⑨ lack economic advantages and are thus left out here. The total reactive power demand of the hydrogen plant under different configurations and typical days (for a single hydrogen price scenario) is shown in Fig. 11. The electrolytic load and reactive power of the four ELZs in a typical renewable generation scenario under configurations ①-③ are illustrated in Figs. 14 to 14, respectively.
@@
As shown in Fig. 14, when the rectifiers in all ELZs are uniform and var resources are adequate, the electrolytic load is evenly distributed across the ELZs to maximize the hydrogen yield, aligning with the equimarginal principle and resulting in [6]. In contrast, mixed rectifier configurations ② and ③, as shown in Figs. 14 and 14, result in ELZs powered by TRs bearing a slightly higher load than those powered by IGBT-Rs in ② and ③.
@@ IV-D Robust Portfolio Results of the Small-scale System for
IV-E Application in a Large-Scale System
@@
For a deviation factor , the optimal rectifier portfolio consists of 24 ELZs powered by 19 TRs and 5 IGBT-Rs, shown in Fig. 16. Notably, the positioning of TRs and IGBT-Rs within the 4-in-1 ELZs, i.e., whether uniform or mixed, is not optimized here and will be explored in future work.
Table VI compares the total revenue for the optimal rectifier configuration, uniform TR/IGBT-R configurations, and ratios used in existing projects for . Compared with the uniform configurations, the optimal configuration increases total revenue by 0.89% to 2.56%, as well as 0.79%, 1.37%, compared with the 1:1 (32 TRs + 32 IGBT-Rs in the Da’an Project [14]) and 1:2 (12 TRs + 24 IGBT-Rs in the Songyuan Project [13]) configurations. The suboptimal portfolio (18 TRs, 6 IGBT-Rs) achieves nearly identical revenue to the optimal portfolio.
@@
From Tables V and VI and Fig. 17, we can summarize that the optimality criterion for MR configurations can be understood as balancing the reactive power in the hydrogen plant, optimizing the investment in SVG to approximately 0, or fully utilizing the reactive power support capability of IGBT-Rs. Excessive investment in IGBT-Rs leads to idle var resources and reduced profitability. Moreover, for low renewable uncertainty, a portfolio of 18 TRs and 6 IGBT-Rs remains near-optimal. Thus, we conclude that the ideal TR-to-IGBT-R ratio for electrolysis rectifiers is approximately 3:1.
| Configuration |
Comparison | ||||
| 24 TRs (uniform) | 1.2248 | 44 | +0.89% | ||||
| 18 TRs + 6 IGBT-Rs (3:1) | 1.2357 | 0 | 0 | ||||
| 12 TRs + 12 IGBT-Rs (1:1) | 1.2261 | 0 | +0.79% | ||||
| 8 TRs + 16 IGBT-Rs (1:2) | 1.2191 | 0 | +1.37% | ||||
| 24 IGBT-Rs (uniform) | 1.2049 | 0 | +2.56% | ||||
| 19 TRs + 5 IGBT-Rs (optimal) | 1.2358 | 0 |
@@ V Conclusions
1) In the optimal portfolio, the reactive power support capability of the IGBT-Rs must be fully utilized, while the need for var compensation approaches zero.
2) The ideal TR-to-IGBT-R ratio for electrolysis rectifiers in ReP2H systems is approximately 3:1.
@@
Future research will focus on the transient behavior and performance of different electrolysis rectifiers, as well as their interactive mechanisms, to enhance the strength and operational stability of ReP2H systems, particularly under weak or off-grid conditions.

### v2

@@ IV-B Investment Portfolio Results of the Small-scale System for
IV-B1 Investment portfolio results
For the small-scale system, with the intraday renewable power uncertainty set to zero, the rectifier configurations of four ELZs are evaluated. Fig. 12 presents the total revenues of different portfolios of TR and IGBT-R, while Table VI summarizes the planning and operational results. Among the results, cases where IGBT-Rs are allowed to adjust reactive power are denoted ②-⑤, and cases where the PF of the IGBT-R is fixed at 1 are denoted ⑥-⑨ (see the discussion about reactive power characteristics in Section II-C-2).
As shown in Fig. 12 and Table VI, when the IGBT-Rs are allowed to provide reactive power, the optimal configuration is 3 TRs and 1 IGBT-R, which better balances investment and operation. In contrast, if the IGBT-R is fixed at , a uniform configuration of 4 TRs is preferable. The reasons are as follows.
@@ IV-B1 Investment portfolio results
In terms of hydrogen production in configurations ① to ⑨, Table VI shows that each additional IGBT-R reduces the hydrogen yield by approximately 2,500 kg/year, resulting in an estimated decrease in revenue of 4.75 CNY. On the other hand, the annual cost of an IGBT-R is 0.104(220–100) = 12.48 CNY greater than that of a TR considering the discount. Therefore, an IGBT-R investment is justified only if it reduces SVG capacity by at least (12.48 + 4.75)/(0.10435) = 4.73 MVar. In particular, Fig. 12 shows the SVG capacity with varying numbers of IGBT-Rs, along with the critical point at which configuring more IGBT-Rs is beneficial. It is clear that the only case ①② meets the conditions.
Moreover, the planning result for the case with 3 ELZs is also discussed, as shown in Table VI. It configures 3 TRs to enhance hydrogen production and deploys SVG to meet the reactive power demand. Compared with the optimal 4-ELZ configuration, although the investment is reduced, reduced RE utilization and operational changes lead to lower operational revenue, performing worse than the 4-ELZ setup.
| ① | 4 + 0 |
6.72 | 1362.64 | 1233.1 | ||||||||||||
| ② | 3 + 1 |
0.55 | 1360.27 | 1229.3 | ||||||||||||
| ③ | 2 + 2 |
0 | 1358.21 | 1218.0 | ||||||||||||
| ④ | 1 + 3 |
0 | 1355.46 | 1217.8 | ||||||||||||
| ⑤ | 0 + 4 |
0 | 1352.77 | 1211.8 | ||||||||||||
| ⑥ | 3 + 1 | 2.0679 | 3.72 | 1360.77 | 1238.3 | ||||||||||||
| ⑦ | 2 + 2 | 2.0618 | 0.56 | 1358.09 | 1235.5 | ||||||||||||
| ⑧ | 1 + 3 | 2.0462 | 0 | 1355.32 | 1228.0 | ||||||||||||
| ⑨ | 0 + 4 | 2.0288 | 0 | 1352.68 | 1217.9 | ||||||||||||
| 3 + 0 (3 ELZs) | 2.0578 (-0.2685+2.3263) | 5.27 | 1288.41 | 1440.3 |
@@
IV-B2 Comparisons with other planning methods and existing configurations
To verify the effectiveness of incorporating active-reactive power coordination and the coupling between hydrogen production and the electrical network into the planning process, and considering the absence of such integrated methods in existing literature, the proposed method is compared against the following intuitive schemes:
M2: Considers the reactive load but does not account for reactive power exchange with the network; instead, the SVG capacity is sized based on the peak reactive demand of the hydrogen plant.s
M3/M4: The number of ELZs is fixed to match an actual project [47] (), while other planning settings for rectifier configuration are kept the same as in M1/M2, respectively.
The planning results are summarized in Table VII. Compared with M1-M4, the proposed method increases total revenue by up to 13.78%. In all M1-M4 cases, SVG capacity is not jointly optimized, and TRs are selected exclusively due to their higher efficiency. However, in M1 and M3, no SVG is deployed to compensate for the reactive demand of TRs. To maintain reactive power balance, the system must rely on alternative sources (e.g., WT, PV, BES) to supply reactive power, leading to excessive var flow in the network. This not only reduces the available capacity for active power delivery, but also exacerbates voltage security issues and increases network losses, ultimately lowering RE utilization and system-level economic performance. In contrast, M2 and M4 neglect the reactive power flexibility of the hydrogen plant, resulting in oversized SVG capacity and poor cost-effectiveness. These results collectively demonstrate that coordinated planning of active and reactive power flows is critical to improving the techno-economic performance of ReP2H systems and should therefore be considered essential.
| Method | Planning Results |
| M1 | 4 TRs + 0 MVar SVG | 1.8267 | +13.78% | ||||
| M2 | 4 TRs + 14.36 MVar SVG | 2.0468 | +1.55% | ||||
| M3 | 3 TRs + 0 MVar SVG | 1.8471 | +12.53% | ||||
| M4 | 3 TRs + 10.77 MVar SVG | 2.0375 | + 2.01% | ||||
| Proposed | Configuration ② | 2.0785 |
@@ IV-C Operational Analysis of Different Rectifier Configurations
IV-C1 Scheduling strategy of ELZs
We analyze power balance, ELZ state transitions, load allocation, and temperature control to demonstrate the core logic and strategy of active power scheduling for hydrogen production. The operational results for a specific scenario using the optimal configuration ② are illustrated in Fig. 13.
As shown in Fig. 13(a), hydrogen production closely follows the fluctuations in renewable generation, with the BES providing load balancing and enabling continuous ELZ operation. The hydrogen plant achieves wide-range power regulation through dynamic switching among production, standby, and idle states across the ELZs. Notably, the 4th ELZ operates at a lower load due to the lower efficiency of its rectifier, while the other three ELZs adhere to the equimarginal principle [6, 16] to maximize hydrogen yield. In addition, thermal management plays a vital role. At h, the 4th ELZ operates at the highest power to use electrolytic heat for P2H efficiency improvement. For details on network-side operations, interested readers are referred to [6], and the reactive power analysis will be discussed later.
IV-C2 Operational comparison
To analyze the operational differences of the ReP2H system under various rectifier configurations, we focus on configurations ①-③. Configurations ④ and ⑤ are similar to ③, whereas ⑥-⑨ lack economic advantages and are thus left out here. The total reactive power demand of the hydrogen plant under different configurations and typical days (for a single hydrogen price scenario) is shown in Fig. 14. The electrolytic load and reactive power of the four ELZs in a typical renewable generation scenario under configurations ①-③ are illustrated in Figs. 17 to 17, respectively.
@@ IV-C2 Operational comparison
As shown in Fig. 17, when the rectifiers in all ELZs are uniform and var resources are adequate, the electrolytic load is evenly distributed across the ELZs to maximize the hydrogen yield, aligning with the equimarginal principle and resulting in [6]. In contrast, mixed rectifier configurations ② and ③, as shown in Figs. 17 and 17, result in ELZs powered by TRs bearing a slightly higher load than those powered by IGBT-Rs in ② and ③.
@@ IV-D Robust Portfolio Results of the Small-scale System for
IV-E Sensitivity Analysis
IV-E1 BES capacity
As the BES participates in both active and reactive power scheduling, it is necessary to examine whether its capacity influences system operation and, consequently, the rectifier portfolio. To this end, a sensitivity analysis is conducted to assess how BES capacity affects the optimal rectifier configuration. The corresponding system revenue is presented in Fig. 19(a). The results show that system revenue increases with BES capacity. However, the relative performance of rectifier configurations remains consistent: , with the MR configuration being the most favorable, in line with the baseline result. This indicates that within a reasonable capacity range, the BES primarily influences active power balancing, while its impact on reactive power flow or the optimal rectifier configuration is negligible due to location differences.
IV-E2 Stack U-I characteristics
Due to manufacturing variability and stack degradation, the performance of electrolysis stacks can vary, primarily reflected in their U-I characteristics. To evaluate the impact of such deviations, we examine rectifier planning under different stack performances. At rated power, cell voltage ranges from 2.0 to 2.2 V (with 2.1 V as the baseline), where lower voltage indicates higher efficiency. Fig. 19(b) shows system revenue under various stack performance levels. It is evident that configuration ② consistently yields the highest revenue. Thus, although stack performance affects overall efficiency, its impact on the relative performance of rectifier configurations is uniform and does not alter the optimal configuration ratio.
IV-E3 SVG cost
To further identify factors influencing rectifier selection, we analyze the effect of SVG cost. Fig. 19(c) presents the variation in system revenue with respect to SVG cost. As SVG becomes cheaper, the revenue associated with the 4-TR configuration increases noticeably, while configurations involving IGBT-Rs are only marginally affected. When SVG cost drops to approximately 250 CNY/kVar, the 4-TR setup becomes the most cost-effective option.
In summary, small variations in most system parameters do not affect the optimal MR configuration, as the self-balancing capability of reactive power within the hydrogen plant ensures an already efficient state. However, SVG cost plays a significant role in economic performance while maintaining a reactive power balance, making it a critical factor in rectifier configuration planning.
IV-F Application in a Large-Scale System
@@
For a deviation factor , the optimal rectifier portfolio consists of 24 ELZs powered by 19 TRs and 5 IGBT-Rs, shown in Fig. 20. Notably, the positioning of TRs and IGBT-Rs within the 4-in-1 ELZs, i.e., whether uniform or mixed, is not optimized here and will be explored in future work. In terms of computational performance, the simulation of a large-scale case is performed on a desktop with an Intel Core i5-12400@2.5 GHz CPU and 16 GB of RAM, and solved by Gurobi 11.0. The convergence of the PH algorithm applied to the SP model is shown in Fig. 21, with a computation time of approximately 24 minutes. This indicates that the large-scale planning problem in this study can be efficiently solved after decomposition, demonstrating the scalability.
Table VIII compares the total revenue for the optimal rectifier configuration, uniform TR/IGBT-R configurations, ratios used in existing projects, and configurations planned by M1 and M2 (introduced in Section IV-B2) for . Compared with the uniform configurations, the optimal configuration increases total revenue by 0.89% to 2.56%, as well as 0.79%, 1.37%, compared with the 1:1 (32 TRs + 32 IGBT-Rs in the Songyuan Project [8]) and 1:2 (12 TRs + 24 IGBT-Rs in the Da’an Project [9]) configurations. Most importantly, the proposed method increases total revenue by at most 10.70% compared to the methods M1 and M2 used in the existing hydrogen planning studies without considering the reactive power and the network. We also find that the suboptimal portfolio (18 TRs, 6 IGBT-Rs) achieves nearly identical revenue to the optimal portfolio.
| Configuration |
| 24 TRs (uniform) | 1.2248 | 44 | +0.89% | |||||
| 18 TRs + 6 IGBT-Rs (3:1) | 1.2357 | 0 | 0 | |||||
| 12 TRs + 12 IGBT-Rs (1:1) | 1.2261 | 0 | +0.79% | |||||
| 8 TRs + 16 IGBT-Rs (1:2) | 1.2191 | 0 | +1.37% | |||||
| 24 IGBT-Rs (uniform) | 1.2049 | 0 | +2.56% | |||||
| 24 TRs (M1) | 1.1164 | 0 | +10.70% | |||||
| 24 TRs (M2) | 1.2113 | 86.16 | +2.02% | |||||
| 19 TRs + 5 IGBT-Rs (optimal) | 1.2358 | 0 |
@@
From Tables VI and VIII and Fig. 22, we can summarize that the optimality criterion for MR configurations can be understood as balancing the reactive power in the hydrogen plant, optimizing the investment in SVG to approximately 0, or fully utilizing the reactive power support capability of IGBT-Rs. Excessive investment in IGBT-Rs leads to idle var resources and reduced profitability. Moreover, for low renewable uncertainty, a portfolio of 18 TRs and 6 IGBT-Rs remains near-optimal. Thus, we conclude that the ideal TR-to-IGBT-R ratio for electrolysis rectifiers is approximately 3:1.
@@ V Conclusions
1) In the optimal portfolio, the reactive power support capability of the IGBT-Rs must be fully utilized to achieve self-balancing of reactive power within the hydrogen plant, while the need for var compensation approaches zero.
2) The ideal TR-to-IGBT-R ratio for electrolysis rectifiers in ReP2H systems is approximately 3:1 given the current costs of rectifiers and SVG.
@@
Future research will focus on the transient behavior and performance of different electrolysis rectifiers, as well as their interactive mechanisms with the electrical network, to enhance the strength and operational stability of ReP2H systems both in steady-state operation and ride-through events, particularly under weak or off-grid conditions.